Information related to Novak Company for 2019 is summarized below.Information related to Novak Company for 2019 is summarized below. Total credit sales Accounts receivable at December 31 Bad
Answer #1

Answer

a)

$ 34900

In direct write-off method the bad debt written off will be the bad debt expense.

b)

Account receivable= $ 853000

5% of 853000= $ 42650

In direct write-off method we use balance in the allowance account. since the account have already a credit balance of $ 2700 , then the amount of bad debt expense

= 42650-2700= $ 39950

c)

5% of $ 853000= $ 42650

Allowance for doubtful account has a debit balance of $ 2700,so amount of bad debt expense,

= 42650+2700= $ 45350